OpenStudy (anonymous):

(3,7) and (8,-1) Find the slope, if it exists, of the line containing the pair of points

OpenStudy (ipwnbunnies):

\[slope = \frac{Y2-Y1}{X2-X1} = \frac{rise}{run}\] Where the coordinates are expressed as (X,Y)

OpenStudy (anonymous):

thank you, this blows my mind :)

OpenStudy (ipwnbunnies):

Lol, no prob.

OpenStudy (anonymous):

ok, it says "the slope M=?"

OpenStudy (ipwnbunnies):

Write, slope is usually associated with the letter 'm'.

OpenStudy (ipwnbunnies):

Right* o.o lol.

OpenStudy (ipwnbunnies):

Just follow the formula.

OpenStudy (anonymous):

ok, I think I am getting it :/

OpenStudy (anonymous):

ok I put it in and it said the answer was -8/5 i may have told you wrong....thank you :)

OpenStudy (ipwnbunnies):

You got it right, good job. :)